OpenStudy (anonymous):

Use summation notation to express the sum 2-1/2+1/8...+1/2048

OpenStudy (anonymous):

looks like you are multiplying each term by \(\frac{1}{4}\) to get the next

OpenStudy (anonymous):

also the signs alternate maybe try \[\sum_{k=0}^n2\times (-\frac{1}{4})^k\]

OpenStudy (anonymous):

oh i guess you have to stop at \(\frac{1}{2048}\) right?

OpenStudy (anonymous):

you could try \[\sum_{k=0}^62\times (-\frac{1}{4})^k\]

OpenStudy (anonymous):

i am sure there are other ways to do it

OpenStudy (anonymous):

Okay thank you